# Randomized Physics-Informed Machine Learning for Uncertainty Quantification in High-Dimensional Inverse Problems

### 11 Dec 2023 v1

@@ 4.3 Convergence of rPICKLE and HMC with the ensemble size
Figure 8 shows that for the low-dimensional case, the convergence properties of HMC and rPICKLE are very similar. We also find that in both methods, the required number of samples for mean and various to reach asymptotic values increases with $\sigma_{r}$. Furthermore, we see that in rPICKLE, the required number of samples is not significantly affected by the dimensionality, which is to be expected as the rPICKLE samples are generated independently from one another.
@@
Figure 8: Convergence of rPICKLE and HMC – $R_{\overline{\xi}_{i}^{b}}(m)$ and $R_{\sigma_{\xi_{i}}^{b}}(m)$ as functions of the ensemble size $m$ for the low- and high-dimensional cases and $\sigma_{r}^{2}=10^{-4}$ and $\sigma_{r}^{2}=10^{-2}$: (a) $\xi_{1}$ of $y_{\mathrm{LD}}$, (b) $\xi_{10}$ of $y_{\mathrm{LD}}$, (c) $\xi_{1}$ of $y_{\mathrm{HD}}$, and (d) $\xi_{100}$ of $y_{\mathrm{LD}}$. HMC results are not available for the high-dimensional case.
@@ 4.4 HMC performance for low- and high-dimensional problems
In this section, we investigate the effect of $\sigma^{2}_{r}$ and the problem dimensionality on the efficiency of HMC.
As mentioned earlier, for certain values of $\sigma^{2}_{r}$, the HMC does not reach the stopping criterion ($10^{4}$ samples) after more than 30 days of running the code. For comparison, rPICKLE generates $10^{4}$ samples in four to five days depending on the value of $\sigma^{2}_{r}$. For the low-dimensional case, the rPICKLE takes approximately 4 minutes to generate the same number of samples for all considered $\sigma^{2}_{r}$ while the computational time of HMC varies from 3 to 6 hours depending on the $\sigma^{2}_{r}$ value.
We find that the increase in the HMC computational time is mainly due to smaller time steps in the integration of the Hamiltonian dynamics equations required to maintain the desirable acceptance rate in the dual averaging algorithm. It was shown in Girolami and Calderhead (2011); Langmore et al. (2023) that the large condition number of the posterior covariance matrix of unknown parameters leads to a decrease in the HMC performance. Here, we demonstrate that the condition number increases with increasing problem dimensionality ($N_{\xi}+N_{\eta}$) and decreasing $\sigma_{r}^{2}$.
@@
where $\bm{\zeta}^{*}$ is the MAP, and $\nabla\nabla\log P(\bm{\zeta}|\mathcal{D}_{res})|_{\bm{\zeta}=\bm{\zeta}^{*}}$ is the Hessian, which we compute by automatic differentiation, evaluated at the MAP. The first-order term vanishes because the gradient of $\log P$ at the MAP is zero. The right-hand side of Eq. (108) is equivalent, up to a constant, to the log probability density of a Gaussian distribution. Under the Laplace assumption, the posterior distribution can be approximated with a Gaussian distribution, which mean is given by the MAP and covariance is found by the inverse of the Hessian evaluated at the MAP point $(\nabla\nabla\log P(\bm{\zeta}|\mathcal{D}_{res})|_{\bm{\zeta}=\bm{\zeta}^{*}})^{-1}$.
Figure 9 shows the eigenvalues (arranged in the descending order) of the approximated posterior covariance for the low- and high-dimensional cases with $\sigma^{2}_{r}=10^{-2}$ and $10^{-4}$. The red dashed lines indicate eigenvalues of the prior covariance (all eigenvalues are equal to one because of the diagonal form of the prior covariance and unit prior variances of the parameters). In the low-dimensional problem, the condition numbers are approximately 21 for $\sigma^{2}_{r}=10^{-2}$ and 1760 for $\sigma^{2}_{r}=10^{-4}$. In the high-dimensional problem, the condition numbers are $\approx 2\times 10^{7}$ for $\sigma^{2}_{r}=10^{-2}$ and $2\times 10^{9}$ for $\sigma^{2}_{r}=10^{-4}$.
@@
Figure 9: Covariance spectrum corresponding to the Laplace approximated posterior at the MAP corresponding to (a) $y_{\mathrm{LD}}$ and (b) $y_{\mathrm{HD}}$ for $\sigma^{2}_{r}=10^{-4}$ and $10^{-2}$. The red dashed line represents the covariance structure informed by the normal prior.
@@ 5 Discussions and Conclusion
The goal of the Bayesian parameter estimation is to obtain a posterior distribution providing a good statistical description of the observed system behavior. Here, we quantify the predictive capacity of the posterior in terms of the distance from the posterior mean or mode to the reference field (could be validation data if the reference field is not available), the percent of coverage, and LPP. This metric can be used for selecting $\sigma^{2}_{r}$ as to minimize the distance and maximize LPP.
We demonstrated that PICKLE provides the mode of the posterior distribution if the PICKLE’s regularization coefficient $\gamma$ is set to $\sigma^{2}_{r}$. Since the PICKLE mode estimation does not require sampling of the posterior, choosing $\sigma^{2}_{r}=\gamma$ to minimize the distance from the posterior mode to the validation data provides a computational advantage. In general, there is no guarantee that $\sigma^{2}_{r}$ selected according to this criterion would also maximize LPP and coverage. However, we found that for considered problems, $\sigma^{2}_{r}$ that minimizes the distance between the mode of the posterior distribution also maximizes the LPP of the posterior distribution. We also found that LPP is more sensitive to $\sigma^{2}_{r}$ than the distance between the mode and data, and, therefore, LPP should be also considered in selecting $\sigma^{2}_{r}$.
@@ Appendix A Gaussian process Regression and the Conditional Covariance
For representing the random field $\mathbf{u}(\mathbf{x};\tilde{\omega})$ conditioned on $u$ measurements, we start by employing Monte Carlo (MC) simulations for computing the prior statistics for $u$. The reason for not directly using another parameterized covariance kernel is because (1) the random field for $u$ is generally not stationary, and (2) the parameterized prior does not enforce the physical constraint. We randomly draw an ensemble of $N_{mc}$ CKLE coefficients $\{\xi_{i}\}_{i=1}^{N_{mc}}$ from $\mathcal{N}(\mathbf{0},\mathbf{I})$, and plug them into Eqn (3) to get $N_{mc}$ realizations of parameters. Then we run numerical simulations to get an ensemble of $N_{mc}$ state predictions. The prior statistics for $u$ can be computed based on the empirical mean and covariance of the ensemble:
@@
where all matrices are counterparts in Eq. (109). We solve the eigenvalue problem for the conditional covariance kernel $\mathbf{C}_{u}^{c}$, to obtain the CKLE representation of the random $u$ field:
@@
ition:relative; bottom:2.2pt;">ATExml[LOGO] body>

### 23 Dec 2023 v2

@@ 4.3 Convergence of rPICKLE and HMC with the Ensemble Size
Figure 8 shows that for the low-dimensional case, the convergence properties of HMC and rPICKLE are very similar. We also find that in both methods, the required number of samples for mean and variance to reach asymptotic values increases with $\sigma_{r}$. Furthermore, we see that in rPICKLE, the required number of samples is not significantly affected by the dimensionality, which is to be expected because the rPICKLE samples are generated independently from one another.
@@
Figure 8: $R_{\overline{\xi}_{i}^{b}}(m)$ and $R_{\sigma_{\xi_{i}}^{b}}(m)$ as functions of the ensemble size $m$ for the low- and high-dimensional cases and $\sigma_{r}^{2}=10^{-4}$ and $\sigma_{r}^{2}=10^{-2}$: (a) $\xi_{1}$ of $y_{\mathrm{LD}}$, (b) $\xi_{10}$ of $y_{\mathrm{LD}}$, (c) $\xi_{1}$ of $y_{\mathrm{HD}}$, and (d) $\xi_{100}$ of $y_{\mathrm{LD}}$. HMC results are not available for the high-dimensional case.
@@ 4.4 HMC Performance for low- and High-dimensional Problems
In this section, we investigate the effect of $\sigma^{2}_{r}$ and OF problem dimensionality on the efficiency of HMC.
As mentioned earlier, for certain values of $\sigma^{2}_{r}$, the HMC does not reach the stopping criterion ($10^{4}$ samples) after more than 30 days of running the code. For comparison, rPICKLE generates $10^{4}$ samples in four to five days depending on the value of $\sigma^{2}_{r}$. For the low-dimensional case, rPICKLE takes approximately 4 minutes to generate the same number of samples for all considered $\sigma^{2}_{r}$, while the computational time of HMC varies from 3 to 6 hours depending on the $\sigma^{2}_{r}$ value.
We find that the increase in HMC computational time is mainly due to smaller time steps in the integration of Hamiltonian dynamics equations required to maintain a desirable acceptance rate in the dual averaging algorithm. It was shown in Girolami and Calderhead (2011); Langmore et al. (2023) that the large condition number of the posterior covariance matrix of unknown parameters leads to a decrease in the HMC performance. Here, we demonstrate that the condition number increases with increasing problem dimensionality ($N_{\xi}+N_{\eta}$) and decreasing $\sigma_{r}^{2}$.
@@
where $\bm{\zeta}^{*}$ is the MAP, and $\nabla\nabla\log P(\bm{\zeta}|\mathcal{D}_{res})|_{\bm{\zeta}=\bm{\zeta}^{*}}$ is the Hessian, which we compute by automatic differentiation, evaluated at the MAP. The first-order term vanishes because the gradient of $\log P$ at the MAP is zero. The right-hand side of Eq. (108) is equivalent, up to a constant, to the log probability density of a Gaussian distribution. Under the Laplace assumption, the posterior distribution can be approximated with a Gaussian distribution, the mean of which is given by the MAP. The covariance is found by the inverse of the Hessian evaluated at the MAP point $(\nabla\nabla\log P(\bm{\zeta}|\mathcal{D}_{res})|_{\bm{\zeta}=\bm{\zeta}^{*}})^{-1}$.
Figure 9 shows the eigenvalues (arranged in descending order) of the approximated posterior covariance for the low- and high-dimensional cases with $\sigma^{2}_{r}=10^{-2}$ and $10^{-4}$. The red dashed lines indicate eigenvalues of the prior covariance (all eigenvalues are equal to one because of the diagonal form of the prior covariance and unit prior variances of the parameters). In the low-dimensional problem, the condition numbers are approximately 21 for $\sigma^{2}_{r}=10^{-2}$ and 1760 for $\sigma^{2}_{r}=10^{-4}$. In the high-dimensional problem, the condition numbers are $\approx 2\times 10^{7}$ for $\sigma^{2}_{r}=10^{-2}$ and $2\times 10^{9}$ for $\sigma^{2}_{r}=10^{-4}$.
@@
Figure 9: Covariance spectrum corresponding to the Laplace approximated posterior for (a) $y_{\mathrm{LD}}$ and (b) $y_{\mathrm{HD}}$ for $\sigma^{2}_{r}=10^{-4}$ and $10^{-2}$. The red dashed line represents the covariance structure informed by the normal prior.
@@ 5 Discussions and Conclusion
The goal of the Bayesian parameter estimation is to obtain a posterior distribution providing a good statistical description of the observed system behavior. Here, we quantify the predictive capacity of the posterior in terms of the distance from the posterior mean or mode to the reference field (could be validation data if the reference field is not available), the percent of coverage, and the LPP. This metric can be used for selecting $\sigma^{2}_{r}$ to minimize the distance and maximize LPP.
We demonstrated that PICKLE provides the mode of the posterior distribution if the PICKLE’s regularization coefficient $\gamma$ is set to $\sigma^{2}_{r}$. Because the PICKLE mode estimation does not require sampling of the posterior, choosing $\sigma^{2}_{r}=\gamma$ to minimize the distance from the posterior mode to the validation data provides a computational advantage. In general, there is no guarantee that $\sigma^{2}_{r}$ selected according to this criterion would also maximize LPP and coverage. However, we found that for considered problems, a $\sigma^{2}_{r}$ that minimizes the distance between the mode of the posterior distribution also maximizes the LPP of the posterior distribution. We also found that LPP is more sensitive to $\sigma^{2}_{r}$ than the distance between the mode and data. Therefore, LPP should also be considered in selecting $\sigma^{2}_{r}$.
@@ Appendix A Gaussian Process Regression and the Conditional Covariance
For representing the random field $\mathbf{u}(\mathbf{x};\tilde{\omega})$ conditioned on $u$ measurements, we start by employing Monte Carlo (MC) simulations for computing the prior statistics for $u$. The reason for not directly using another parameterized covariance kernel is because (1) the random field for $u$ is generally not stationary, and (2) the parameterized prior does not enforce the physical constraint. We randomly draw an ensemble of $N_{mc}$ CKLE coefficients $\{\xi_{i}\}_{i=1}^{N_{mc}}$ from $\mathcal{N}(\mathbf{0},\mathbf{I})$, and plug them into Eq. (3) to get $N_{mc}$ realizations of parameters. Then, we run numerical simulations to get an ensemble of $N_{mc}$ state predictions. The prior statistics for $u$ can be computed based on the empirical mean and covariance of the ensemble:
@@
where all matrices are counterparts in Eq. (109). We solve the eigenvalue problem for the conditional covariance kernel $\mathbf{C}_{u}^{c}$ to obtain the CKLE representation of the random $u$ field: